OpenStudy (jennychan12):

Explain this? A skier traveling 12.0 m/s reaches the foot of a steady upward 18° incline and glides 12.2 m up along this slope before coming to rest. What is the average coefficient of friction?

OpenStudy (jennychan12):

i know that 

OpenStudy (jennychan12):

i know it's non conservative, but what about the mass? because the W(fr) has no mass attached to it.

OpenStudy (anonymous):

Find the average retardation by v^2-u^2=2as where final velocity v=0 initial velocity u=12m/s distance travelled=12.2m After that use a=gsin(18)+xgcos(18) to solve for x, the coefficent of friction

OpenStudy (jennychan12):

so the masses basically just cancel out? thanks.